Determine which of the following lines, if any, are perpendicular or parallel
Line A: 3y+x=6; line b: y=3x+2; line c: 3x+y=0

Answer by ewatrrr(24785)

 
Hi


A: 3y+x=6 OR y = (-1/3)x+2 |Slope m = (-1/3)
B: y=3x+2 |Slope m = 3/1
C: 3x+y=0 OR y = -3x |Slope m = -3
Line A and B represent perpendicular lines
(slopes are negative reciprocals of one another, that is (-1/3)*3/1 = -1